# Modelling Quantum Transduction for Multipartite Entanglement Distribution

### v2

@@ Abstract
Index Terms:
Quantum Internet, Quantum Transduction, Entanglement Distribution, Multipartite Entanglement, Teleporting, Microwave, Optical
@@ II Problem statement: Direct vs Teleported Multipartite Entanglement Distribution
Figure 2: TMD: Teleported Multipartite entanglement Distribution for the same multipartite state considered in Fig. 1. First, the orchestrator generates and share three EPR pairs with the clients, which represent the communication resource utilized for teleporting the multipartite state to the clients. Then, by performing local operations – i.e., Bell State Measurements (BSMs) – between the ebits of the EPR pairs at the orchestrator and the ebits of the GHZ state, followed by classical communications [13], the $3$-qubit GHZ is shared between 3 clients.
@@
In the considered scenario, the orchestrator is a superconducting node, performing quantum computations on multipartite entangled qubits. The rational for this design setup choice lies in its killer application, namely the distributed quantum computation [17] in a quantum data center (also known in literature as quantum server farm) [18]. Indeed, this constitutes a very hot research topic, which is gathering a huge attention by technology giants such as Amazon [19], IBM [20] and Cisco [21, 22, 23]. This attention is not surprising, since, in a short-term time-horizon, interconnecting different quantum processors with a quantum local-area network (Q-LAN) represents the only way (with current technology-readiness level) to scale the number of qubits beyond the thousands that are necessary to achieve fault-tolerant quantum computing [16, 17]. Recently we showed in [16, 24, 25, 26] that, by leveraging the local operations at the orchestrator, it is possible to boost the network performance in terms of overhead and reliability with respect to the scenario without a supernode.
In this context, distributing multipartite entangled states among remote superconducting quantum nodes constraints the adopted distribution strategy. Indeed, the chosen strategy must take into account the huge frequency gap between intra- and inter-nodes frequencies – namely, between microwave and optical frequencies – by resorting to quantum transduction. And, we can distinguish two main categories for multipartite entanglement distribution strategies:
@@
It is worthwhile to mention that, in the figures, we considered a $3$-qubit GHZ [^1] A Greenberger–Horne–Zeilinger (GHZ) state – formally, |$\ket{GHZ}=\frac{1}{\sqrt{2}}\left(\ket{0}^{\otimes n}+\ket{1}^{\otimes n}\right)$ for a $n$-qubit state [28] – is a maximally entangled state, characterised by maximally connectivity. Indeed, a state is maximally connected if, for any two qubits, there exists a sequence of single-qubit measurements on the remaining qubits that, when performed, guarantee that the two qubits end up in a maximally entangled state [29]. On the other hand, GHZ states exhibit minimum persistency equal to $1$. Indeed, the persistency of a multipartite entangled state is the minimum number of qubits that need to be measured to guarantee that the resulting state is unentangled [29, 3]. state to be distributed to the clients. This choice is made only for the sake of pictorial representation, since our modelling is not limited to this class of multipartite entanglement or to the number of clients. Indeed, the proposed strategies hold for a generic $n$-qubit multipartite state to be distributed to $n$ clients [^2] To be more precise, the proposed analysis and strategies work also for the general case where the number of client is $m\leq n$. This results into a multipartite entangled state shared between the orchestrator and clients, i.e., the orchestrator retains some qubit at its side as in [16, 24, 25, 26]. However, in the following, for the sake of pictorial representation, we consider the scenario where the number of ebit of the multipartite state coincides with the number of clients..
@@ III-A Quantum Transduction Efficiency
Regarding extraction ratios $\zeta_{x}$, values of $\zeta_{x}$ close to $1$ map into internal cavity losses $\gamma_{x,i}\stackrel{{\scriptstyle\triangle}}{{=}}\gamma_{x}-\gamma_{x,e}$ being relatively small [33]. It mush be noted that there is a wide-scientific consensus in considering unitary values for $\zeta_{x}$ feasible to achieve in the near-future. And indeed, typical values assumed in theoretical studies are around $\zeta_{x}=0.9$ [32], whereas experimental values in the order of $0.1-0.2$ have already been measured [48]. On the contrary, unitary cooperativity is still considered beyond the state-of-the art in the near-future, with experimental values for $C$ measured recently reaching around $0.3$ [34]. In the following subsections, we study and discuss the impact of these hardware constraints on the key functionality represented by entanglement distribution.
@@ III-B TMD with “vanilla” quantum transduction
As mentioned at the beginning of this section, in TMD with vanilla quantum transduction [^8] In the following, when there is no ambiguity issue, we could refer to TMD strategy with vanilla quantum transduction as vanilla TMD., one ebit of each EPR – locally generated at the orchestrator – is distributed to the clients with a sequence of up- and down-conversion. Thus, we can model the TMD strategy with vanilla quantum transduction as in Fig. 2. Specifically, $n$ microwave EPR pairs |$\ket{\Phi^{o,o}_{M,M}}$ are generated at the orchestrator, where the superscript |$\ket{\cdot^{o}}$ denotes the orchestrator as “location” of each ebit of the EPR and the subscript |$\ket{\cdot_{M}}$ denotes the microwave photon domain of each ebit. Once the generation is performed, one ebit of each EPR pair generated is up-converted to optical domain, transmitted to each client and down-converted to the microwave domain therein. The process eventually results in an EPR state, denoted as |$\ket{\Phi^{o,c}_{M,M}}$, since it is shared between the orchestrator $o$ and the arbitrary client $c\in\{1,\ldots,n\}$. Accordingly, once all the $n$ EPRs are distributed through the network, the overall state |$\ket{\Omega^{v}}$ obtained with vanilla TMD before teleportation is given by:
|$\ket{\Omega^{v}}=\ket{\Pi}\otimes\ket{\Phi^{o,c}_{M,M}}^{\otimes n},$ (6)
with |$\ket{\Pi}$ denoting the multipartite entangled state to be distributed to the $n$ clients. The teleportation of the multipartite entangled state can be now be performed.
@@
with $l_{o,c}$ denoting the length of the fiber link between orchestrator and client $c$, $L_{0}$ denoting the attenuation length of the fiber [^9] As for today, commercial fibers feature an attenuation lower than $1$db/km. As instance, optical photons with wavelength equal to $1550$nm – i.e., DWDM ITU $100$GHz channel number $35$ in the C band – experience an attenuation of $0.2$dB/km, which corresponds to $L_{0}=22$ km [50]. and $\eta^{o}_{\uparrow}$ and $\eta^{c}_{\downarrow}$ denoting the efficiency of the up- and down-conversion at the orchestrator and at the client, respectively.
@@ III-C TMD with intrinsic-entanglement quantum transduction (IE-TMD)
|$\ket{\Phi^{o,o}_{M,O}}=c_{0}\ket{0_{M}^{o}0_{O}^{o}}+c_{1}\ket{1_{M}^{o}1_{O}^{o}}.$ (8)
with the subscripts $(\cdot_{M})$ and $(\cdot_{O})$ denoting the photon domain – i.e., microwave or optical – and the superscripts $(\cdot^{o})$ and $(\cdot^{c})$ denoting the “location” of each ebit, i.e., at the orchestrator or at the client. Accordingly, in (8) the term |$\ket{1_{M}^{o}1_{O}^{o}}$ denotes the event in which SPDC successfully generated a couple of photons, one at microwave and the other at optical frequency. Conversely, the term |$\ket{0_{M}^{o}0_{O}^{o}}$ denotes the event in which SPDC failed, and no microwave nor optical photon have been generated. The coefficients $c_{0}$ and $c_{1}$ depends on the transducer hardware parameters. Entanglement in the form of (8) is generated whenever the quantum transducer is initialized with no input – namely, no photon to be converted as in Fig. 5 – and the input pump field is set to operate on a frequency that is the sum of the frequencies of the optical and microwave photons (aka “blue detuning”).
@@
Two assumptions underlying equation (8) must be better discussed. Regarding the assumption of generating a two-level state – i.e., the assumption of restricting the admissible system state to 2-level Fock states for each field rather than an uncountable levels – it is not restrictive, since it simply requires a specific inizialization of the microwave field inside the cavity as in Fig. 5 [57]. In this case, an input pump field is set to operate on a frequency that is the difference of the frequencies of the optical and microwave photons (aka “red detuning”). The pump is exploited to obtain an entangled state in the form of |$\ket{\Psi^{o,o}_{M,O}}=c_{0}\ket{0^{o}_{M}1^{o}_{O}}+c_{1}\ket{1^{o}_{M}0^{o}_{O}}$ , which is equivalent to (8) up to a basis change [57, 61].
The term |$\ket{0_{M}^{o}1_{O}^{o}}$ denotes a successful conversion of the initialized microwave field within the transducer, while the term |$\ket{1_{M}^{o}0_{O}^{o}}$ denotes a failed conversion. Again, the coefficients $c_{0}$ and $c_{1}$ depends on the transducer features. Depending on a careful setting of the transduction hardware parameters, it is possible to obtain an EPR state (see Appendix C) in the form:
|$\ket{\Psi^{o,o}_{M,O}}=\frac{1}{\sqrt{2}}(\ket{0^{o}_{M}1^{o}_{O}}+\ket{1^{o}_{M}0^{o}_{O}})$ (9)
@@
|$\ket{\Phi^{o,c}_{M,M}}=\frac{1}{\sqrt{2}}(\ket{0_{M}^{o}0_{M}^{c}}+\ket{1_{M}^{o}1_{M}^{c}})$ (10)
Accordingly, once all the $n$ EPR pairs are distributed through the network, the overall state |$\ket{\Omega^{IE}}$ obtained with IE-TMD is equivalent to the state given in (6), and the teleportation of the multipartite entangled state can be now be performed.
@@ III-D TMD with intrinsic-entanglement quantum transduction & swapping (IES-TMD)
Accordingly, the overall state |$\ket{\Omega^{IES}}$, after the swapping-like distribution procedure is once again equivalent to the state given in (6), and the teleportation of the multipartite entangled state can be performed.
@@
It is important to highlight that, when both transducers generate optical photons, only one detector click is triggered due to path erasure. However, in this case, the state shared between the orchestrator and the client is |$\ket{0^{o}_{M}}\ket{0^{c}_{M}}$, which is definitely not an entangled state as in (14). However, if we reasonably assume the availability of photon-number-resolved detectors (PNRD), then it is possible to distinguish the event of receiving two optical photons – one for each transducer in each link – from the event where only one optical photon is received. And the double-photon event can be discarded in favour of a new distribution attempt.
@@ IV Discussion
Stemming from the developed analysis, it appears that the nature of the entangled pairs exploited in the TMD strategies are diverse. More into details, in the case of Vanilla-TMD, the EPR pairs, generated by the orchestrator, are at microwave frequencies. Therefore, in order to be distributed to the clients, one ebit for each EPR pair has to be up-converted to optical frequency through the transducer available at the orchestrator. On the contrary, in the case of IE-TMD and IES-TMD strategies, hybrid entangled pairs are generated: one ebit of each EPR pair is at microwave frequency and the other ebit is at optical frequency. Thus, by adopting IE-TMD or IES-TMD strategy, the EPR generation process is directly performed at the transducer hardware, by exploiting two mode squeezing or beam splitter interaction, respectively, as detailed in Sec. III-C. The aforementioned difference in the type and nature of the exploited EPR pairs in the different TMD strategies are summarized in Tab.II.
We graphically summarise in Fig. 11 the communication performances of the proposed strategies, in terms of achievable quantum capacity, as a function of the cooperativity $C$.
It is worthwhile to stress that the transducer efficiency does not provide sufficient granularity to grasp all the mechanisms/phenomena underlying a transducer process, as the type of encoding implemented within the transducer. Thus, we theoretically evaluated upper bound for the ebit distribution probabilities and thus for the quantum capacities. This is the reason for which in the legend of Fig. 11 we indicated “capacity upper bounds”. For the sake of clarity, Tab. III shows the strategies comparison in terms of achievable quantum capacity for the required values of cooperativity $C$.
@@
The aforementioned consideration allows us to further stress an aspect highlighted in the final remark of Sec. III-A. Specifically, as observed, quantum transduction is not just a merely frequency conversion process. Indeed, several factors impact the transducer conversion efficiency and consequently the failure conversion probability [11, 70]. Nevertheless, the hardware literature still lacks of a general model able to grasp all the mechanisms/phenomena underlying the electro-optical transducer process from the entanglement generation and distribution perspective. Stemming from this, and by considering that our main goal is to start a dialogue among different communities (i.e. communication, physics, and hardware), we choose to propose a first model of quantum transducer in which the conversion efficiency constitutes the main parameter for characterizing its performances. The rational for this choice is that several factors determining the conversion failure can be embedded within the conversion efficiency. An example to all, to ensure a successful conversion, the mode phase-matching condition has to be fulfilled. The effect of mode mismatching is typically mitigated through filtering techniques [71], that reduces the probability of conversion and consequently the conversion efficiency. As result, the mode-mismatch can be embedded in the conversion efficiency parameter through the parameter $g$ in (1), as shown in [11]. Our design choice allows not only to abstract from the particulars of the specific technology underlying the transducer hardware, but also to easily and progressively take into account technological advancements that improve the transducer performances.
@@ Appendix A Proof of Lemma 2
(QEC), where the incoming ebit is replaced by the erasure state |$\ket{\epsilon}$ with probability $(1-p^{v}_{c})$. Formally, the output density matrix $\rho_{o}$ on a given orchestrator-client link is given by:
$\rho_{o}=p^{v}_{c}\rho_{i}+\left(1-p^{v}_{c}\right)\ket{\epsilon}\bra{\epsilon},$ (17)
@@ Appendix C Proof of Lemma 4
where $T+R=1$, $RT^{*}+TR^{*}=0$ and with $T=\eta^{o}_{\uparrow}$ and $R=1-\eta^{o}_{\uparrow}$. In terms of Fock states [76], given an input state |$\ket{\psi_{in}}=b^{\dagger}_{in}\ket{00}=\ket{1_{M}0_{O}}$ the output state can be expressed as follows:
|$\ket{\psi_{out}}$ $=\sqrt{\eta^{o}_{\uparrow}}\ket{0_{M}1_{O}}+\sqrt{1-\eta^{o}_{\uparrow}}\ket{1_{M}0_{O}}.$ (25)
Whenever $\eta^{o}_{\uparrow}=T=1$, |$\ket{\psi_{in}}$ is up-converted into |$\ket{\psi_{out}}=\ket{0_{M}1_{O}}$ with $100\%$ up-conversion efficiency and the hardware behave as an ideal quantum transducer. Conversely, if $\eta^{o}_{\uparrow}=\frac{1}{2}$, the $L$ matrix becomes [77]:
@@
and, in this condition the transducer acts as a 50/50 lossless beam-splitter [77] generating a Bell State |$\ket{\psi_{out}}=\frac{1}{\sqrt{2}}(\ket{0^{o}_{M}1^{o}_{O}}+\ket{1^{o}_{M}0^{o}_{O}})$. From (25), for any setting of operative parameters resulting in $\eta^{o}_{\uparrow}=\frac{1}{2}$ – i.e., $C=C_{th}=3-2\sqrt{2}$ – an EPR pair is generated. Conversely, whenever $C\neq C_{th}$, we have two cases. For degenerate values of $\eta^{o}_{\uparrow}$ – i.e., $\eta^{o}_{\uparrow}=0$ or $1$ – the transducer outputs a single mode and no entanglement is generated, whereas for non-degenerated values of $\eta^{o}_{\uparrow}$ an odd superposition of states |$\ket{01}$ and |$\ket{10}$ – which is not a maximally entangled state – is generated.

### v4

@@ II Problem statement: Direct vs Teleported Multipartite Entanglement Distribution
Figure 2: TMD: Teleported Multipartite entanglement Distribution for the same multipartite state considered in Fig. 1. First, the orchestrator generates and share three EPR pairs with the clients, which represent the communication resource utilized for teleporting the multipartite state to the clients. Then, by performing local operations – i.e., Bell State Measurements (BSMs) – between the ebits of the EPR pairs at the orchestrator and the ebits of the GHZ state, followed by classical communications [13], the 33-qubit GHZ is shared between 3 clients.
@@
In the considered scenario, the orchestrator is a superconducting node, performing quantum computations on multipartite entangled qubits. The rational for this design setup choice lies in its killer application, namely the distributed quantum computation [17] in a quantum data center (also known in literature as quantum server farm) [18]. Indeed, this constitutes a very hot research topic, which is gathering a huge attention by technology giants such as Amazon [19], IBM [20] and Cisco [21, 22, 23]. This attention is not surprising, since, in a short-term time-horizon, interconnecting different quantum processors with a quantum local-area network (Q-LAN) represents the only way (with current technology-readiness level) to scale the number of qubits beyond the thousands that are necessary to achieve fault-tolerant quantum computing [16, 17]. Recently we showed in [16, 24, 25, 26] that, by leveraging the local operations at the orchestrator, it is possible to boost the network performance in terms of overhead and reliability with respect to the scenario without a supernode. In this context, distributing multipartite entangled states among remote superconducting quantum nodes constraints the adopted distribution strategy. Indeed, the chosen strategy must take into account the huge frequency gap between intra- and inter-nodes frequencies – namely, between microwave and optical frequencies – by resorting to quantum transduction. And, we can distinguish two main categories for multipartite entanglement distribution strategies:
@@
[…]
@@ III-A Quantum Transduction Efficiency
Regarding extraction ratios $\zeta_{x}$, values of $\zeta_{x}$ close to $1$ map into internal cavity losses $\gamma_{x,i}\stackrel{{\scriptstyle\triangle}}{{=}}\gamma_{x}-\gamma_{x,e}$ being relatively small [33]. It mush be noted that there is a wide-scientific consensus in considering unitary values for $\zeta_{x}$ feasible to achieve in the near-future. And indeed, typical values assumed in theoretical studies are around $\zeta_{x}=0.9$ [32], whereas experimental values in the order of 0.10.20.1-0.2 have already been measured [48]. On the contrary, unitary cooperativity is still considered beyond the state-of-the art in the near-future, with experimental values for $C$ measured recently reaching around $0.3$ [34]. In the following subsections, we study and discuss the impact of these hardware constraints on the key functionality represented by entanglement distribution.
@@ III-B TMD with “vanilla” quantum transduction
[…]
@@
with $l_{o,c}$ denoting the length of the fiber link between orchestrator and client $c$, $L_{0}$ denoting the attenuation length of the fiber [^9] As for today, commercial fibers feature an attenuation lower than 11db/km. As instance, optical photons with wavelength equal to 15501550nm – i.e., DWDM ITU 100100GHz channel number $35$ in the C band – experience an attenuation of 0.20.2dB/km, which corresponds to $L_{0}=22$ km [50]. and $\eta^{o}_{\uparrow}$ and $\eta^{c}_{\downarrow}$ denoting the efficiency of the up- and down-conversion at the orchestrator and at the client, respectively.
@@ III-C TMD with intrinsic-entanglement quantum transduction (IE-TMD)
[…]
@@
[…]
@@
[…]
@@ III-D TMD with intrinsic-entanglement quantum transduction & swapping (IES-TMD)
[…]
@@
[…]
@@ IV Discussion
Stemming from the developed analysis, it appears that the nature of the entangled pairs exploited in the TMD strategies are diverse. More into details, in the case of Vanilla-TMD, the EPR pairs, generated by the orchestrator, are at microwave frequencies. Therefore, in order to be distributed to the clients, one ebit for each EPR pair has to be up-converted to optical frequency through the transducer available at the orchestrator. On the contrary, in the case of IE-TMD and IES-TMD strategies, hybrid entangled pairs are generated: one ebit of each EPR pair is at microwave frequency and the other ebit is at optical frequency. Thus, by adopting IE-TMD or IES-TMD strategy, the EPR generation process is directly performed at the transducer hardware, by exploiting two mode squeezing or beam splitter interaction, respectively, as detailed in Sec. III-C. The aforementioned difference in the type and nature of the exploited EPR pairs in the different TMD strategies are summarized in Tab.II.We graphically summarise in Fig. 11 the communication performances of the proposed strategies, in terms of achievable quantum capacity, as a function of the cooperativity $C$. It is worthwhile to stress that the transducer efficiency does not provide sufficient granularity to grasp all the mechanisms/phenomena underlying a transducer process, as the type of encoding implemented within the transducer. Thus, we theoretically evaluated upper bound for the ebit distribution probabilities and thus for the quantum capacities. This is the reason for which in the legend of Fig. 11 we indicated “capacity upper bounds”. For the sake of clarity, Tab. III shows the strategies comparison in terms of achievable quantum capacity for the required values of cooperativity $C$.The aforementioned consideration allows us to further stress an aspect highlighted in the final remark of Sec. III-A. Specifically, as observed, quantum transduction is not just a merely frequency conversion process. Indeed, several factors impact the transducer conversion efficiency and consequently the failure conversion probability [11, 70]. Nevertheless, the hardware literature still lacks of a general model able to grasp all the mechanisms/phenomena underlying the electro-optical transducer process from the entanglement generation and distribution perspective. Stemming from this, and by considering that our main goal is to start a dialogue among different communities (i.e. communication, physics, and hardware), we choose to propose a first model of quantum transducer in which the conversion efficiency constitutes the main parameter for characterizing its performances. The rational for this choice is that several factors determining the conversion failure can be embedded within the conversion efficiency. An example to all, to ensure a successful conversion, the mode phase-matching condition has to be fulfilled. The effect of mode mismatching is typically mitigated through filtering techniques [71], that reduces the probability of conversion and consequently the conversion efficiency. As result, the mode-mismatch can be embedded in the conversion efficiency parameter through the parameter $g$ in (1), as shown in [11]. Our design choice allows not only to abstract from the particulars of the specific technology underlying the transducer hardware, but also to easily and progressively take into account technological advancements that improve the transducer performances.
@@ Appendix A Proof of Lemma 2
[…]
$|,\rho_{o}=p^{v}_{c}\rho_{i}+\left(1-p^{v}_{c}\right)\ket{\epsilon}\bra{\epsilon}$, (17)
@@ Appendix C Proof of Lemma 4
[…]
@@
[…]